OpenStudy (anonymous):

What is the solution to the equation? 2 x squared plus 4 equals 20 Choose all answers that are correct. A. − 2 square root of 2 B. − 2 square root of 8 C. 2 square root of 2 D. 2 square root of 8

OpenStudy (anonymous):

@egbeach

OpenStudy (egbeach):

\[2x ^{2} +4=20\]

OpenStudy (anonymous):

yup

OpenStudy (egbeach):

move the 4 to the other side

OpenStudy (anonymous):

so 2x^2=20-4

OpenStudy (anonymous):

2x^2=16

OpenStudy (egbeach):

how would you get ride of the 2?

OpenStudy (anonymous):

divide both sides

OpenStudy (egbeach):

so do that...

OpenStudy (anonymous):

x^2=8

OpenStudy (egbeach):

and whats the square root of 8?

OpenStudy (egbeach):

\[2\sqrt{2}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it says choose all answers that are correct

OpenStudy (anonymous):

so that's one

OpenStudy (anonymous):

@egbeach

OpenStudy (egbeach):

when you square something what do you always end up with?

OpenStudy (anonymous):

umm..

OpenStudy (anonymous):

so A?

OpenStudy (anonymous):

@egbeach

OpenStudy (egbeach):

should be

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Thanks! :)